OpenStudy (bookworm14):

1 question !!! I got the other ones right, but this one is confusing me please help !!! (will post it below) --->

OpenStudy (bookworm14):

ANY HELP !!!! @satellite73 @hilbertboy96 @TheLukeskywalker2

OpenStudy (anonymous):

\[n+\left( n-2 \right)\times \frac{ 70 }{ 100 }=8.5 \]

OpenStudy (bookworm14):

why would you have 2 (n)s in the problem >

OpenStudy (bookworm14):

and what would be the value of "n"

OpenStudy (bookworm14):

@surjithayer @agent0smith @mathstudent55 @ganeshie8

OpenStudy (anonymous):

is the answer correct?

OpenStudy (bookworm14):

I don't know because i have to write an equation and solve for "X" but i fi get this wrong i have to start all over, and im just really trying to figure it out

OpenStudy (bookworm14):

I thought the equation might be\[(n-2)+\frac{ 70 }{ 100 } =8.5\]

OpenStudy (bookworm14):

in which case n=9.8

OpenStudy (anonymous):

multiply by 100 \[100 n+70 n-70 *2=\frac{ 85 }{ 10 }*100\] 170 n=850+140 170 n=990 \[n=\frac{ 990 }{ 170 }=\frac{ 99 }{ 17 }\] After increasing n becomes 8.5 so n <8.5

OpenStudy (anonymous):

first n is the number second n is for to calculate increase.

OpenStudy (bookworm14):

i got it wrong , here is the explanation they gave me:

OpenStudy (bookworm14):

now i have to start over ;-;

OpenStudy (anonymous):

instead of n they are saying n-2 becomes 8.5 after increase. i thought n becomes 8.5 after increase.

OpenStudy (bookworm14):

hmm do you think you could assist on about 5 more , it started me over with new questions

OpenStudy (anonymous):

sure

OpenStudy (bookworm14):

make that 4, i just got one right (its khan academy i have to get 5 correct in a row to pass the topic) here is the quest, i will try to solve also

OpenStudy (anonymous):

3x+2=50

OpenStudy (anonymous):

find x

OpenStudy (bookworm14):

I got \[\frac{ 50-2 }{ 3 }=x\] but yours looks more accurate x=16 i think

OpenStudy (anonymous):

3x=50-2=48 \[x=\frac{ 48 }{ 3 }=16\]

OpenStudy (bookworm14):

your equation was right ..here is the next one

OpenStudy (anonymous):

\[\frac{ c }{ 2 }+1.2=3.7\]

OpenStudy (bookworm14):

\[(3.7-1.2)*2=c\] why are mine turning out so different from yours?

OpenStudy (bookworm14):

mine is the one solving for c correct? or is it your because it wants us to write an equation solving for C

OpenStudy (bookworm14):

mine was right

OpenStudy (anonymous):

both are correct \[\frac{ c }{ 2 }=3.7-1.2\] \[c=\left( 3.7-1.2 \right)*2\]

OpenStudy (bookworm14):

I know, but mine was worder the way they wanted